OpenStudy (daniellelovee):

Forty-two percent of the members in a book club prefer to read nonfiction books. Use the Table of Random Digits on pages A-3 to A-4 in the reference guide to simulate randomly selecting and asking 15 members, “Do you prefer to read fiction or nonfiction books?” Let the digits 01 through 42 represent a response of nonfiction. (a) Start at the beginning of line 122 and write the first 30 digits as two-digit numbers. Circle those that represent a response of nonfiction. (b) Write p and for the situation.

OpenStudy (daniellelovee):

@ganeshie8

OpenStudy (anonymous):

could i help or just leave cause some people just want a specific person

OpenStudy (daniellelovee):

you can help :)

OpenStudy (anonymous):

are u homeschool

OpenStudy (daniellelovee):

why are you asking?

OpenStudy (anonymous):

cuz never done this question sorry

OpenStudy (anonymous):

before

OpenStudy (daniellelovee):

oh thats ok I'll just wait for @ganeshie8 then thannks for trying

OpenStudy (anonymous):

not can atleast try first

OpenStudy (daniellelovee):

ok

OpenStudy (anonymous):

where are the choices again

OpenStudy (daniellelovee):

there are no choices

OpenStudy (anonymous):

thats right mybad forgot

OpenStudy (anonymous):

sorry nervous

OpenStudy (daniellelovee):

is ok why are you nervous?

OpenStudy (anonymous):

just things and this question at once

OpenStudy (daniellelovee):

oh ok lol

OpenStudy (anonymous):

i think to write the situation.first

OpenStudy (anonymous):

thats what im doing

OpenStudy (daniellelovee):

alright

OpenStudy (anonymous):

90510 mabe

OpenStudy (anonymous):

what do u think it is

OpenStudy (daniellelovee):

I know how to find A but idk how to find B

OpenStudy (anonymous):

i can't see it to

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

imma try again ok

OpenStudy (daniellelovee):



OpenStudy (daniellelovee):

alright

OpenStudy (anonymous):

ok thnxs

OpenStudy (daniellelovee):

np

OpenStudy (anonymous):

Start at the beginning of line 122 and write the first 30 digits as two-digit numbers. Circle those that represent a response of nonfiction.

OpenStudy (anonymous):

should work right for this question

OpenStudy (anonymous):

nvm nervous again

OpenStudy (daniellelovee):

why 2 digits instead of 3?

OpenStudy (anonymous):

thats right you are very smarter than should have thought of that darn it

OpenStudy (daniellelovee):

lol is ok just think it through a bit more and dont get nervous

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

232 is triple digit

OpenStudy (anonymous):

one of them

OpenStudy (anonymous):

right

OpenStudy (daniellelovee):

i believe so

OpenStudy (anonymous):

lets try it

OpenStudy (anonymous):

sorry took to long

OpenStudy (daniellelovee):

is alright im falling asleep either way lol

OpenStudy (anonymous):

alright go ahead i'll probably still be here trying to figure out the answer

OpenStudy (daniellelovee):

alright thank you :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

cardinal

OpenStudy (anonymous):

well what u think it is

OpenStudy (anonymous):

yeah i think its 232 86% positive

OpenStudy (anonymous):

carlo are you here or still typing

OpenStudy (anonymous):

try 232 or 230 or 217

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

back

OpenStudy (anonymous):

So as the prompt states, the first 30 digits of line 122 are supposed to be written in groups of two like so: 47, 71, 18, 92, 00, 20, 40, 89, 02, 26, 48, 33, 77, 87, 84, 74 And those values who fall between 01 and 42 represent a yes to nonfiction. These are: 18, 20, 40, 02, 26, and 33 Now we find the P value. Remember, the null hypothesis is 42% preference to nonfiction. So our sample percentage is: \[\frac{ 6 }{ 30 } \times 100 = 20\] which means from our sample, there's only 20% preference for nonfiction which is significantly lower than our initial hypothesis. So our sample proportion is: \[P = 0.20\] And the standard error of our sample percentage is: \[\sqrt{\frac{ p(1-p) }{ n }} = \sqrt{\frac{ 0.2(1-0.2) }{ 30 }} = \sqrt{\frac{ 0.2(0.8) }{ 30 }}\] we then simplify to: \[E _{s} = 0.07\] now to find the standard score: \[\frac{ proportional - initial }{ error } = \frac{ 0.20-0.42 }{ 0.07 } = -3.1\] This means that our sample score is 3.1 standard errors apart from the initial score. So to compare with a standard z-table, we find that for a standard error of 0.07 and a difference of 3.1 standard errors between the proportional and initial scores, we have a 0.0008 (or 0.08%) probability of Z (20% sample) being less than its value. And therefore, we can conclude that P-value is 0.0016 and also that the 42% hypothesis is highly probable in this situation.

OpenStudy (anonymous):

finally trying to figure out y s long she's slepping right now so me and you buddy

OpenStudy (anonymous):

the answer twenty -20

OpenStudy (anonymous):

how so? @youngestDJ

OpenStudy (anonymous):

from what u typed asked u if the answer is twenty -20 i think 232 how bout u @Cardinal_Carlo

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

I don't know where your 232 is coming from. Please elaborate.

OpenStudy (anonymous):

sorry sticky keys

OpenStudy (anonymous):

what do u think the answer is

OpenStudy (anonymous):

I'd like @Daniellelovee to be the final judge.

OpenStudy (anonymous):

she's sleep me and u are here only